OpenStudy (anonymous):

I need help with setting up this problem James, Kim, and Katrina each bought lemonade at different stores. James paid $6 per gallon, Kim paid $8 per gallon, and Katrina paid $11 per gallon. Between the three of them, 25 gallons were purchased. If Kim bought 4 more gallons than Katrina and they pooled their lemonade together to sell for $8 per gallon, find the number of gallons James, Kim and Katrina each purchased.

OpenStudy (anonymous):

K = Kim J = James X = Katrina K + J + X = 25 K = X + 4 K + J + X = 25 X + 4 + J + X = 25 2X = 21 - J X = (21 - J)/2 this is not the answer....I am stuck....Let me tag someone

OpenStudy (anonymous):

@mathstudent55 ....help

OpenStudy (anonymous):

@whpalmer4 ...help

OpenStudy (anonymous):

@genius12 ...help

OpenStudy (anonymous):

K8 + J6 + X11 = 8 ??? That doesn't make sense

OpenStudy (anonymous):

@satellite73 ...help

OpenStudy (anonymous):

oh....8 * 25

hero (hero):

k = katrina j = james m = kim k + j + m = 25 11k + 6j + 11m = 8*25 m = k + 4

OpenStudy (anonymous):

I had you 2nd equation but I didn't multiply the 8 * 25

OpenStudy (anonymous):

typo...your

OpenStudy (anonymous):

Why does the 8*25 matter?

hero (hero):

"they pooled their lemonade together to sell for $8 per gallon"

OpenStudy (anonymous):

and they bought 25 gal

OpenStudy (anonymous):

So the 8*25=200 and would you plug numbers into the second equation to get the answer?

hero (hero):

You solve the system of two equations in two variables, j and k after making the proper substitutions.

OpenStudy (anonymous):

11k+6j+11m=200 -11 -11 6j+11m=189 -11 -11 6j/6=178/6 j=29.6 repeating can't be right

OpenStudy (anonymous):

you cant subtract 11 if its 11k

OpenStudy (anonymous):

then what would you do

OpenStudy (anonymous):

we know that m = k + 4

hero (hero):

Sorry, it's 8m not 11m

OpenStudy (anonymous):

Would you solve that equation first

hero (hero):

k = katrina j = james m = kim k + j + m = 25 11k + 6j + 8m = 8*25 m = k + 4

hero (hero):

Since I messed up, I might as well just post the steps. I meant 8m the first time...it was a typo.

OpenStudy (anonymous):

thanks Hero...your a....well...a hero....your name speaks for itself :)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Kristen.....medal for Hero

hero (hero):

Actually, there's another typo. I should have gotten 200 for that second equation.

hero (hero):

I plugged the numbers back in but I didn't get 200 for the second equation

hero (hero):

And I know why...

OpenStudy (anonymous):

my eye blinked....I didn't even notice

hero (hero):

k + j + (k + 4) = 25 11k + 6j + 8(k + 4) = 200 2k + j = 21 19k + 6j = 168 12k + 6j = 126 19k + 6j = 168 7k = 42 k = 6 m = k + 4 = 6 + 4 = 10 j = 9

hero (hero):

There we go, that should work.

OpenStudy (anonymous):

good catch :)

hero (hero):

that's what I like about algebra. You can always plug back in to find your mistakes.

OpenStudy (anonymous):

Where did you get 12 and 126?

hero (hero):

I multiplied both sides of 2k + j = 21 by 6

hero (hero):

I know I may have confused you with what I said beforehand.

OpenStudy (anonymous):

I understand now thank you:)